OpenStudy (anonymous):

PLEASE!!!! I NEED THIS ANSWER IN 15 MINUTES!!!! A particle travels 17 times around a 15-cm radius circle in 30 seconds. What is the average speed (in m/s) of the particle?

OpenStudy (anonymous):

What is the circumference of a circle?\[C = 2 \pi R\]This is the distance the object travels in one trip around the circle. Therefore, total distance\[d = 17 \cdot C = 34 \pi R\] Velocity is defined as\[v = {\Delta d \over \Delta t} = {34 \pi R \over 30}\]

OpenStudy (anonymous):

thank you is R the radius?